1/^x2 + 5/x + 6 = 0
Can you help me solve this equation?

Answer by rothauserc(4718)

***********************************
is 1/^x2 really 1/(x^2)?, if yes then
***********************************
1/^x2 + 5/x + 6 = 0
multiply both sides of = by x^2
1 + 5x + 6x^2 = 0
rearrange terms
6x^2 + 5x + 1 = 0
factor polynomial
(3x + 1) * (2x + 1) = 0
the two solutions are
x = -1/3 and x = -1/2
**********************************
check the solutions
1) x = -1/3
(1 / (1/9)) + (5/(-1/3)) + 6 = 0
9 - 15 + 6 = 0
0 = 0
**********************************
2) x = -1/2
(1 / (1/4)) + (5/(-1/2)) + 6 = 0
4 -10 + 6 = 0
0 = 0
our answers check